OpenStudy (pokemon23):

solve for w: P=2L+2W solve for w: P=2L+2W @Mathematics

OpenStudy (saifoo.khan):

P=2L+2W 2W = P - 2L \[W = \frac{P-2L}{2}\]

OpenStudy (pokemon23):

you did it again :0

OpenStudy (saifoo.khan):

Omg!